OpenStudy (the_puppet1):

Solve the equation 4y - 3 = 1 for y. A. -1 B. 0 C. 1 D. 2

OpenStudy (carolina4567):

Answer ?

OpenStudy (the_puppet1):

I think it's C

OpenStudy (carolina4567):

There u go :)

OpenStudy (dannyrod2000):

y=1

OpenStudy (dannyrod2000):

Simplifying 4y + -3 = 1 Reorder the terms: -3 + 4y = 1 Solving -3 + 4y = 1 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '3' to each side of the equation. -3 + 3 + 4y = 1 + 3 Combine like terms: -3 + 3 = 0 0 + 4y = 1 + 3 4y = 1 + 3 Combine like terms: 1 + 3 = 4 4y = 4 Divide each side by '4'. y = 1

OpenStudy (dannyrod2000):

do you understand??